# Zero-Shot EEG-to-Gait Decoding via Phase-Aware Representation Learning

### v1

@@ Abstract
Accurate decoding of lower-limb motion from EEG signals is essential for advancing brain–computer interface (BCI) applications in movement intent recognition and control. However, challenges persist in achieving causal, phase-consistent predictions and in modeling both inter- and intra-subject variability. To address these issues, we propose NeuroDyGait, a domain-generalizable EEG-to-motion decoding framework that leverages structured contrastive representation learning and relational domain modeling. The proposed method employs relative contrastive learning to achieve semantic alignment between EEG and motion embeddings. Furthermore, a multi-cycle gait reconstruction objective is introduced to enforce temporal coherence and maintain biomechanical consistency. To promote inter-session generalization, during fine-tuning, a domain dynamic decoding mechanism adaptively assigns session-specific prediction heads and learns to mix their outputs based on inter-session relationships. NeuroDyGait enables zero-shot motion prediction for unseen individuals without requiring adaptation and achieves superior performance in cross-subject gait decoding on benchmark datasets. Additionally, it demonstrates strong phase-detection capabilities even without explicit phase supervision during training. These findings highlight the potential of relational domain learning in enabling scalable, target-free deployment of BCIs.
{IEEEkeywords}
domain generalization, electroencephalography, contrastive learning, gait decoding, zero-shot prediction
@@ 1 Introduction
Restoring human mobility using brain-computer interfaces (BCIs) is an emerging focus in neurorehabilitation [colucci2022brain]. A key component of this effort is the decoding of lower-limb motor intent from non-invasive electroencephalography (EEG) signals [he2018brain]. This capability is essential for enabling assistive technologies such as exoskeletons and neuroprosthetics [kilicarslan2021neuro]. Although recent progress has demonstrated the feasibility of EEG-based gait decoding, significant challenges remain for real-world deployment. A major limitation lies in the limited availability of normal motion data from patients with neurological impairments. These individuals are often unable to perform standard gait patterns. This makes it critical to design decoders that generalize across individuals without relying on patient-specific fine-tuning, thereby enabling practical and scalable use in rehabilitation settings [azab2019weighted, fahimi2019inter].
@@
Despite this progress, two key challenges hinder reliable decoding of gait dynamics. The first limitation concerns the widespread adoption of segment-to-point prediction strategies. These methods use a short window of historical EEG signals to estimate the motion state at a single final time step [he2018mobile, brantley2018full, shafiul2020prediction, presacco2011neural].This formulation satisfies the requirements of causal decoding. However, it reduces human motion to isolated time points, neglecting the temporal continuity and biomechanical constraints that characterize natural gait. In practice, the context of gait provides critical information for inferring joint angles, as joint configurations are strongly influenced by surrounding temporal dynamics. Human locomotion is intrinsically structured and phase-dependent: even under similar motor intent, joint trajectories can vary significantly across individuals, and identical joint configurations may reflect different gait phases depending on context and prior states [hamacher2017exploring, presacco2012decoding]. Without explicitly modeling this structure, predictions risk becoming unstable or semantically inconsistent. Therefore, a robust decoding framework must incorporate these underlying dynamics to enable the model to differentiate between gait states that are superficially similar but functionally distinct across different cycles.
@@
The second challenge lies in the limitations of current cross-domain learning strategies, which often treat multiple sessions or subjects as independent sources. These methods typically use shared encoders for generalization or subject-specific heads for individual differences [jayaram2016transfer, chen2021ms], but fail to capture structured relationships across domains. This oversimplification becomes particularly problematic when generalizing across heterogeneous data sources, where domain shifts—caused by factors such as sensor placement, individual differences, and recording conditions—introduce structured dependencies between domains. Empirical evidence indicates that both inter-session and inter-subject variability can substantially degrade the consistency and reliability of biosignal decoding. This highlights the importance of models that not only capture domain-specific characteristics but also exploit shared structure across domains to enable more robust and transferable representation learning [huang2023discrepancy, maswanganyi2022statistical, ng2023deep]. Therefore, a robust framework should be designed to model both intra-domain specificity and inter-domain structure to enable more adaptive and generalizable learning.
@@ 2.1 EEG-Based Motor Execution Decoding
EEG-based neural decoding of lower-limb motor execution has been widely studied, leveraging EEG’s high temporal resolution for capturing motor control dynamics. Pfurtscheller and colleagues [pfurtscheller2001motor] demonstrated that event-related desynchronization (ERD) and synchronization (ERS) effectively characterize motor execution and imagery, laying the groundwork for decoding lower-limb movements. Ang and colleagues [ang2012filter] used common spatial pattern (CSP) and linear discriminant analysis (LDA) to decode gait phases, enhancing motor-related EEG feature extraction. Schirrmeister et al. [schirrmeister2017deep] employed convolutional neural networks (CNNs), boosting accuracy in classifying lower-limb motor imagery. Recent deep learning methods have further improved spatiotemporal feature extraction from EEG signals [goh2018spatio, tortora2020deep, fu2022matn], demonstrating their ability to capture complex nonstationary neural patterns. Finally, Wang et al. [wang2018implementation] developed a real-time BCI system for lower-limb exoskeleton control, validating the feasibility of EEG-driven motor decoding in rehabilitation settings. While these works significantly advance EEG-based BCIs, they typically focus on classifying discrete movement states or rely on pointwise regression, often lacking biomechanical awareness and dynamic consistency in continuous motion decoding.
@@ 2.2 Cross-Subject Generalization in EEG Decoding
A major challenge in EEG decoding lies in inter-subject variability, where differences in brain dynamics, head geometry, and sensor placement cause distributional shifts between individuals. To address this, domain adaptation methods have been applied to align source and target feature distributions through adversarial training [li2018cross], statistical moment matching [long2015learning], or subspace projection [liang2022multi]. However, most approaches assume access to target-domain data during training, which is often unrealistic in clinical settings where normative data from impaired users are unavailable. Domain generalization methods aim to overcome this by learning domain-invariant representations solely from source data [wang2024dmmr, lu2023hybrid], yet often treat all source domains independently or equally. In practice, EEG and motor signals vary both across and within subjects, and ignoring inter-session relationships can limit generalization.
In this work, we model session-level structure explicitly by assigning each source session a dedicated decoder and learning a dynamic mixture of their outputs. This relation-aware strategy captures both intra-session specificity and cross-session similarity, improving robustness in zero-shot subject transfer.
@@ 2.3 Contrastive Representation Learning for Structured Movement
In video-based action recognition, motion-aware frameworks such as MaCLR align visual and motion modalities for enhanced video representations [xiao2022maclr], while MCL emphasizes motion through alignment of gradient maps and optical flow [li2021motion]. For skeleton-based action recognition, contrastive learning captures structural dynamics [guo2022contrastive], with HiCLR enforcing consistency across hierarchical augmentations [zhang2023hierarchical], and cross-modality approaches modeling complex patterns [li2023cross].
In EEG-based motor imagery, contrastive learning addresses inter-subject variability using CNNs and attention mechanisms [li2024self], while supervised contrastive learning improves gait recognition using EEG and EMG signals [fu2023gait].
@@ 3.1 Overview
Our proposed framework, NeuroDyGait, consists of two training stages: (1) a pretraining stage (see Stage I in Fig.1) that learns temporally structured and semantically meaningful EEG embeddings by reconstructing synchronized motion signals and modeling cross-domain relationships, optimized using a combination of reconstruction, contrastive, and prediction losses; and (2) a domain generalization stage (see Stage II in Fig.1) that employs a session-wise head architecture with a learnable domain fusion mechanism, optimized via a domain fusion loss and a supervised prediction loss.
Figure 1: Overview of NeuroDyGait architecture. Stage I: The dual-encoder model consists of an EEG encoder, a motor encoder, and a decoder. It is trained with reconstruction, contrastive, and prediction losses to extract biomechanically meaningful EEG representations. Stage II: During training, session-specific heads and a domain weighting layer are optimized to predict final motor state. Test: The model computes a normalized mixture of all source-domain heads for zero-shot generalization.
@@ 3.2.2 Reconstruction Loss
$\mathcal{L}_{\text{rec}}=\frac{1}{N}\sum_{i=1}^{N}\|\hat{\mathbf{y}}_{i}-\mathbf{y}_{i}\|^{2}$
@@ 3.2.3 Prediction Loss
$\mathcal{L}_{\text{pred}}=\frac{1}{N}\sum_{i=1}^{N}\|\hat{\mathbf{y}}_{i,T}-\mathbf{y}_{i,T}\|^{2}$
where $\hat{\mathbf{y}}_{i,T}$ and $\mathbf{y}_{i,T}$ denote the predicted and ground truth motor states at the final time step $T$ for the $i$-th sample, and $N$ is the number of training samples in a batch.
@@ 3.2.4 Relative Contrastive Learning
Inspired by Xu et al.  [xu2024relcon], we introduce a relative contrastive learning framework that captures fine-grained semantic relationships between EEG and motor signals. Unlike traditional contrastive methods that define fixed positive and negative sets, our approach treats all intra-batch samples as candidates and enforces relative similarity rankings between them. This encourages the model to learn an embedding space where the distance structure reflects the semantic alignment between modalities.
@@
Given an EEG embedding $\mathbf{z}_{e}\in\mathbb{R}^{d}$ and a candidate motor embedding $\mathbf{z}_{m}\in\mathbb{R}^{d}$, we define a cross-attention mechanism that treats $\mathbf{z}_{e}$ as the query and $\mathbf{z}_{m}$ as the key and value. The attention weight is computed via a scaled dot-product:
$\alpha=\text{softmax}\left(\frac{\langle W_{q}\mathbf{z}_{e},W_{k}\mathbf{z}_{m}\rangle}{\sqrt{d}}\right)$ (1)
@@
$\hat{\mathbf{z}}_{\text{m}}=W_{o}(\alpha\cdot W_{v}\mathbf{z}_{m})$ (2)
where $W_{o}\in\mathbb{R}^{d\times d}$ are learnable projection matrices. $W_{o}$ maps the aggregated attention output back to the shared latent space. This allows the attended motor representation $\hat{\mathbf{z}}_{\text{m}}$ to be directly compared to the EEG embedding $\mathbf{z}_{\text{e}}$ for distance computation, enabling cross-modal alignment in a unified embedding space.
The cross-modal distance is measured as the mean squared error between the EEG embedding and the attended motor representation:
$d(\mathbf{z}_{e},\mathbf{z}_{m})=\left|\hat{\mathbf{z}}_{m}-\mathbf{z}_{e}\right|_{2}^{2}$ (3)
This learned distance function serves as a semantic similarity measure between EEG and motor embeddings, capturing alignment based on shared motion-related features.
@@
Using the pairwise distances $d(\mathbf{z}_{e}^{(i)},\mathbf{z}_{m}^{(j)})$ between all EEG-motor pairs in a batch, we define a relative contrastive loss that enforces ranking consistency. For each anchor $i$, we compute the similarity matrix:
@@
where $\tau$ is a temperature scaling parameter, and $b$ is a learnable scalar bias. For each anchor, we sort all candidates by similarity and compute a listwise loss based on log-softmax over the sorted similarities:
$\mathcal{L}_{i}=-\sum_{j=1}^{B-1}\left(S_{ij}-\log\sum_{k=j}^{B-1}e^{S_{ik}}\right)$ (5)
Here, we exclude the self-similarity term $S_{ii}$ to avoid trivial comparisons, resulting in a total of $B-1$ candidates for each anchor in the loss computation.
The final relative contrastive loss is averaged across all anchors:
$\mathcal{L}_{\text{rcl}}=\frac{1}{B}\sum_{i=1}^{B}\mathcal{L}_{i}$ (6)
This loss encourages the model to assign higher similarity to more semantically aligned motor sequences while preserving the relative ranking among all candidates.
The total loss of stage I is a weighted sum of all components:
$\mathcal{L}_{\text{stage I}}=\mathcal{L}_{\text{rec}}+\lambda_{\text{pred}}\mathcal{L}_{\text{pred}}+\lambda_{\text{rcl}}\mathcal{L}_{\text{rcl}}$
We set $\lambda_{\text{pred}}=0.5$ to moderately encourage accurate motor prediction during pretraining without dominating the learning dynamics, allowing the model to prioritize representation alignment through the relative contrastive loss. Meanwhile, $\lambda_{\text{rcl}}=1$ emphasizes the preservation of semantic consistency in EEG-motor relations, which is crucial for robust downstream decoding.

### v2

# EEG-to-Gait Decoding via Phase-Aware Representation Learning

@@ Abstract
Accurate decoding of lower-limb motion from EEG signals is essential for advancing brain-computer interface (BCI) applications in movement intent recognition and control. This study presents NeuroDyGait, a two-stage, phase-aware EEG-to-gait decoding framework that explicitly models temporal continuity and domain relationships. To address challenges of causal, phase-consistent prediction and cross-subject variability, Stage I learns semantically aligned EEG–motion embeddings via relative contrastive learning with a cross-attention-based metric, while Stage II performs domain relation–aware decoding through dynamic fusion of session-specific heads. Comprehensive experiments on two benchmark datasets (GED and FMD) show substantial gains over baselines, including a recent 2025 model EEG2GAIT. The framework generalizes to unseen subjects and maintains inference latency below 5 ms per window, satisfying real-time BCI requirements. Visualization of learned attention and phase-specific cortical saliency maps further reveals interpretable neural correlates of gait phases. Future extensions will target rehabilitation populations and multimodal integration.
keywords:
domain generalization, electroencephalography, contrastive learning, gait decoding
journal: Neural Networks
@@ 1 Introduction
Restoring human mobility using brain-computer interfaces (BCIs) is an emerging focus in neurorehabilitation [6]. A key component of this effort is the decoding of lower-limb motor intent from non-invasive electroencephalography (EEG) signals [15]. This capability is essential for enabling assistive technologies such as exoskeletons and neuroprosthetics [21]. Although recent progress has demonstrated the feasibility of EEG-based gait decoding, significant challenges remain for real-world deployment. A major limitation lies in the limited availability of normal motion data from patients with neurological impairments. These individuals are often unable to perform standard gait patterns. This makes it critical to design decoders that generalize across individuals without relying on patient-specific fine-tuning, thereby enabling practical and scalable use in rehabilitation settings [3, 8].
@@
Despite this progress, two key challenges hinder reliable decoding of gait dynamics. The first limitation concerns the widespread adoption of segment-to-point prediction strategies. These methods use a short window of historical EEG signals to estimate the motion state at a single final time step [16, 4, 38, 36].This formulation satisfies the requirements of causal decoding. However, it reduces human motion to isolated time points, neglecting the temporal continuity and biomechanical constraints that characterize natural gait. In practice, the context of gait provides critical information for inferring joint angles, as joint configurations are strongly influenced by surrounding temporal dynamics. Human locomotion is intrinsically structured and phase-dependent: even under similar motor intent, joint trajectories can vary significantly across individuals, and identical joint configurations may reflect different gait phases depending on context and prior states [14, 35]. Without explicitly modeling this structure, predictions risk becoming unstable or semantically inconsistent. Therefore, a robust decoding framework must incorporate these underlying dynamics to enable the model to differentiate between gait states that are superficially similar but functionally distinct across different cycles.
@@
The second challenge lies in the limitations of current cross-domain learning strategies, which often treat multiple sessions or subjects as independent sources. These methods typically use shared encoders for generalization or subject-specific heads for individual differences [19, 5], but fail to capture structured relationships across domains. This oversimplification becomes particularly problematic when generalizing across heterogeneous data sources, where domain shifts—caused by factors such as sensor placement, individual differences, and recording conditions—introduce structured dependencies between domains. Empirical evidence indicates that both inter-session and inter-subject variability can substantially degrade the consistency and reliability of biosignal decoding. This highlights the importance of models that not only capture domain-specific characteristics but also exploit shared structure across domains to enable more robust and transferable representation learning [17, 31, 33]. Therefore, a robust framework should be designed to model both intra-domain specificity and inter-domain structure to enable more adaptive and generalizable learning.
@@
In summary, NeuroDyGait adopts a two-stage EEG-to-gait decoding framework: Stage I learns phase-aware, semantically structured EEG–motion embeddings through multi-objective pretraining, and Stage II performs relation-aware decoding by dynamically fusing session-specific heads based on learned inter-domain attention. This design enables the model to capture both temporal gait dynamics and structured cross-domain relationships for robust cross-subject prediction. We have released the full implementation as open-source at https://github.com/FuXi1999/NeuroDyGait.
@@ 2.1 EEG-Based Motor Execution Decoding
EEG-based neural decoding of lower-limb motor execution has been widely studied, leveraging EEG’s high temporal resolution for capturing motor control dynamics. Pfurtscheller and colleagues [34] demonstrated that event-related desynchronization (ERD) and synchronization (ERS) effectively characterize motor execution and imagery, laying the groundwork for decoding lower-limb movements. Ang and colleagues [2] used common spatial pattern (CSP) and linear discriminant analysis (LDA) to decode gait phases, enhancing motor-related EEG feature extraction. Schirrmeister et al. [37] employed convolutional neural networks (CNNs), boosting accuracy in classifying lower-limb motor imagery. Recent deep learning methods have further improved spatiotemporal feature extraction from EEG signals [12, 42, 11], demonstrating their ability to capture complex nonstationary neural patterns. Finally, Wang et al. [44] developed a real-time BCI system for lower-limb exoskeleton control, validating the feasibility of EEG-driven motor decoding in rehabilitation settings. While these works significantly advance EEG-based BCIs, they typically focus on classifying discrete movement states or rely on pointwise regression, often lacking biomechanical awareness and dynamic consistency in continuous motion decoding.
@@ 2.2 Cross-Subject Generalization in EEG Decoding
A major challenge in EEG decoding lies in inter-subject variability, where differences in brain dynamics, head geometry, and sensor placement cause distributional shifts between individuals. To address this, domain adaptation methods have been applied to align source and target feature distributions through adversarial training [24], statistical moment matching [29], or subspace projection [28]. However, most approaches assume access to target-domain data during training, which is often unrealistic in clinical settings where normative data from impaired users are unavailable. Domain generalization methods aim to overcome this by learning domain-invariant representations solely from source data [45, 30], yet often treat all source domains independently or equally. In practice, EEG and motor signals vary both across and within subjects, and ignoring inter-session relationships can limit generalization.
In this work, we model session-level structure explicitly by assigning each source session a dedicated decoder and learning a dynamic mixture of their outputs. This relation-aware strategy captures both intra-session specificity and cross-session similarity, improving robustness in unseen-subject transfer.
@@ 2.3 Contrastive Representation Learning for Structured Movement
In video-based action recognition, motion-aware frameworks such as MaCLR align visual and motion modalities for enhanced video representations [47], while MCL emphasizes motion through alignment of gradient maps and optical flow [26]. For skeleton-based action recognition, contrastive learning captures structural dynamics [13], with HiCLR enforcing consistency across hierarchical augmentations [52], and cross-modality approaches modeling complex patterns [23].
In EEG-based motor imagery, contrastive learning addresses inter-subject variability using CNNs and attention mechanisms [27], while supervised contrastive learning improves gait recognition using EEG and EMG signals [9].
@@ 3 Methods
To improve clarity and reproducibility, Table 1 summarizes all symbols and parameters used in the framework.
Symbol Definition
$\mathbf{x}\in\mathbb{R}^{C\times T}$ EEG signal segment (C channels, T time points)
$\mathbf{y}\in\mathbb{R}^{J\times T}$ Motion (joint angle) sequence (J joints, T frames)
$f_{e}(\cdot)$, $f_{m}(\cdot)$ EEG encoder and motor encoder networks
$g(\cdot)$ Motor decoder network
$\mathbf{z}_{e}$, $\mathbf{z}_{m}$ EEG and motor latent embeddings
$\hat{\mathbf{y}}$, $\hat{\mathbf{y}}_{T}$ Reconstructed motion sequence and final predicted frame
NN Number of samples per mini-batch
$d$ Latent embedding dimension
$W_{q},W_{k},W_{v},W_{o}\in\mathbb{R}^{d\times d}$ Query, key, value, and output projection matrices in cross-attention
$\eta$ Cross-attention coefficient (Stage I)
$\tau$ Temperature parameter in contrastive learning
$b$ Learnable bias for similarity scaling
$d(\cdot,\cdot)$ Cross-modal distance measure between EEG and motor embeddings
$\mathcal{L}_{\text{rec}}$, $\mathcal{L}_{\text{pred}}$, $\mathcal{L}_{\text{rcl}}$ Reconstruction, prediction, and relative contrastive losses
$h_{s}(\cdot)$ Session-specific prediction head for domain/session $s$
$\mathcal{L}_{\text{sup}}$, $\mathcal{L}_{\text{df}}$ Supervised and domain-fusion losses in Stage II
$\boldsymbol{\alpha}\in[0,1]^{N_{\text{src}}}$ Domain weighting vector predicted by scoring network (Stage II)
$W_{a}$, $\mathbf{b}_{a}$ Parameters of the domain scoring network
$\mathbf{m}\in\{0,1\}^{N_{\text{src}}}$ One-hot mask excluding current session
$\hat{\mathbf{y}}_{\text{mix}}$, $\hat{\mathbf{y}}_{\text{test}}$ Mixture predictions during training and inference
Table 1: Summary of main notations used in the proposed framework.
@@ 3.1 Overview
Our proposed framework, NeuroDyGait, consists of two training stages: (1) a pretraining stage (Stage I) that learns temporally structured and semantically meaningful EEG embeddings by reconstructing synchronized motion signals and modeling cross-domain relationships, optimized using a combination of reconstruction, contrastive, and prediction losses; and (2) a domain generalization stage (Stage II) that employs a session-wise head architecture with a learnable domain fusion mechanism, optimized via a domain fusion loss and a supervised prediction loss.
Figure 1: Overview of NeuroDyGait architecture. Stage I: The dual-encoder model consists of an EEG encoder, a motor encoder, and a decoder. It is trained with reconstruction, contrastive, and prediction losses to extract biomechanically meaningful EEG representations. Stage II: During training, session-specific heads and a domain weighting layer are optimized to predict final motor state. Test: The model computes a normalized mixture of all source-domain heads for unseen-domain generalization.
@@ 3.2.2 Reconstruction Loss
$\mathcal{L}_{\text{rec}}=\frac{1}{N}\sum_{i=1}^{N}\|\hat{\mathbf{y}}_{i}-\mathbf{y}_{i}\|^{2}$ (1)
@@ 3.2.3 Prediction Loss
$\mathcal{L}_{\text{pred}}=\frac{1}{N}\sum_{i=1}^{N}\|\hat{\mathbf{y}}_{i,T}-\mathbf{y}_{i,T}\|^{2}$ (2)
where $\hat{\mathbf{y}}_{i,T}$ and $\mathbf{y}_{i,T}$ denote the predicted and ground truth motor states at the final time step $T$ for the $i$-th sample.
@@ 3.2.4 Relative Contrastive Learning
Inspired by Xu et al. [48], we introduce a relative contrastive learning framework that captures fine-grained semantic relationships between EEG and motor signals. Unlike traditional contrastive methods that define fixed positive and negative sets, our approach treats all intra-batch samples as candidates and enforces relative similarity rankings between them. This encourages the model to learn an embedding space where the distance structure reflects the semantic alignment between modalities.
@@
Given an EEG embedding $\mathbf{z}_{e}\in\mathbb{R}^{d}$ and a candidate motor embedding $\mathbf{z}_{m}\in\mathbb{R}^{d}$, we define a cross-attention mechanism that treats $\mathbf{z}_{e}$ as the query and $\mathbf{z}_{m}$ as the key and value. The attention coefficient is computed via scaled dot-product:
$\color[rgb]{0,0,0}\eta=\text{softmax}\!\left(\frac{\langle W_{q}\mathbf{z}_{e},W_{k}\mathbf{z}_{m}\rangle}{\sqrt{d}}\right)$} (3)
@@
$\hat{\mathbf{z}}_{\text{m}}=W_{o}({\color[rgb]{0,0,0}\eta}\cdot W_{v}\mathbf{z}_{m})$ (4)
The cross-modal distance is then defined as:
$d(\mathbf{z}_{e},\mathbf{z}_{m})=\|\hat{\mathbf{z}}_{m}-\mathbf{z}_{m}\|_{2}^{2}$ (5)
@@
Using pairwise distances $d(\mathbf{z}_{e}^{(i)},\mathbf{z}_{m}^{(j)})$ between all EEG–motor pairs in a batch, we define a relative contrastive loss enforcing ranking consistency:
@@
For each anchor, we compute the listwise loss:
$\mathcal{L}_{i}=-\sum_{j=1}^{N-1}\left(S_{ij}-\log\sum_{k=j}^{N-1}e^{S_{ik}}\right)$ (7)
and the total contrastive loss:
$\mathcal{L}_{\text{rcl}}=\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}_{i}$ (8)
The final Stage I objective is:
$\mathcal{L}_{\text{stage I}}=\mathcal{L}_{\text{rec}}+\mathcal{L}_{\text{pred}}+\mathcal{L}_{\text{rcl}}$ (9)
Loss Interaction Rationale
All Stage I losses are equally weighted, as this configuration achieves stable convergence and balanced gradients without explicit tuning. Combining sequence reconstruction and final-frame prediction encourages the encoder to capture temporal continuity beneficial for Stage II fine-tuning rather than overfitting to instantaneous frames.